OpenStudy (anonymous):

Can someone please help me? I just want to check my answer. I'm supposed to convert y + 2 = - ( x - 2 ) into standard form Ax + By = C I have the answer but I just want to double check it to be sure. Thanks! :)

OpenStudy (anonymous):

My answer is x + y = 0 I don't know if it's correct or not. Help?

OpenStudy (adamk):

That's incorrect. First, let's distribute the negative inside those parenthesis. What would you get?

OpenStudy (adamk):

Also, I think you're missing an equal sign.

OpenStudy (anonymous):

Whoops. y + 2 = - ( x - 2 ) y + 2 = -x + 2

OpenStudy (adamk):

Ah, that's better. Ok, good you distributed correctly. Then add "x" to both sides. y + x + 2 = 2 Then, subtract 2 x + y = 0. Good you got it right! Just wasn't sure about that first equation.

OpenStudy (anonymous):

Oh ok. Thanks, I wasn't sure though because I had never gotten that before. :/ Okay well thanks for the help. :)